OpenStudy (anonymous):

What is the value of p – (–3.5) when p = 8.2? Thanks if you help!(I'm horrible at math!)

Vocaloid (vocaloid):

p = 8.2 means that you can replace p with 8.2 so, p - (-3.5) = 8.2 - (-3.5)

jimthompson5910 (jim_thompson5910):

also, the two negatives cancel to make a positive example: 2 - (-3) = 2+3 = 5